(6) Determine the general solution x(a) to a dx 2do = a(1 - ax).
Answer #1

game = a 11-ax) 2 da dal = 20 (1-9x) da X da = 2-2an => dat 2ax = 2 da In first order linear one of yt Py =q form) pea) = 2a